Tuesday, December 4, 2012
More work on the 147.39 repeater...
We retuned all of the receive & transmit cavities, and found them all to be out of adjustment...some by quite a bit. They were last adjusted back in 2002. We changed the configuration of the cavities to improve isolation, and have an additional pass cavity on the receive and transmit ports. De-sense is gone, and all is well with that aspect of the repeater, now.
All 4 cooling fans were replaced. They run non-stop, and have turned many revolutions since the installation of this repeater at the Clayton site in 2002. All of the heat sinks are cleaned, and the repeater is running nice and cool.
A temporary repair had been made a while back with a jumper lead, and that was removed and hard wired to correct a control logic problem. There is one more similar issue, but I forgot my glasses and couldn't see the terminal that needed to be soldered :-)
Weak signal tests were made, and it is responding to signals that are right down in the noise, so we're pretty happy with that. A few programming changes were made, and that wrapped up the work for the day.
The "to-do" list is getting short now with only a few items left like: replacing a defective PL encoder board on the transmitter, repairing a control logic issue for the 2nd voice announcer module, and replacing a backup battery on the voice IDers.
Next, the work on the Clayton site begins...